the market price is in a perfectly competitive market is 100$. a perfectly competitive firm has a marginal cost given by MC=2+2Q. Find the profit maximizing output and revenue for this firm.


1
Expert's answer
2021-05-30T15:43:45-0400
"Solution"

Given that :

"market \ price,P=\$100\\
MC=2+2Q"


Profit maximization requires that :"P=MC"

"100=2+2Q\\
2Q=98\\
Q=\frac{98}{2}=49"

The firms revenue will be :"TR=P\times Q=100\times 49=\$4,900"